OpenStudy (anonymous):

12 y^3/8 x^4* 6x^6/20 y^8

OpenStudy (anonymous):

\[9x^2/20y^5\], i think, if i've interpreted your question correctly :P

OpenStudy (callisto):

\[\frac{12 y^3}{8 x^4}\times \frac{ 6x^6}{20 y^8}\]\[=\frac{12\times6 x^6y^3}{8 \times 20x^4y^8}\]\[=\frac{9x^{6-4}}{20y^{8-3}}\]\[ =\frac{9x^{2}}{20y^{5}}\]

OpenStudy (anonymous):

callisto, how do you get those fractions, all straight and all??? Mine are all WHOOPS.

OpenStudy (callisto):

Hmm??? \frac{numerator}{denominator}

OpenStudy (anonymous):

oh thats cool thanks :) lemme try dat out: \[\frac{i like}{rainbows}\] WOAHHH, THATS MAJORLY COOL.

OpenStudy (callisto):

try : i \ like to add a 'space' between i and like

OpenStudy (unklerhaukus):

right click on some one else's text and you can show as math tex commands and copy the script

OpenStudy (callisto):

:O ^Smart!

OpenStudy (callisto):

\frac{i \ like}{rainbows}

OpenStudy (anonymous):

You can't just do that, you have to type it IN the equations thingy :P

OpenStudy (unklerhaukus):

\[\left.\frac{\text{ spectrum }}{\mathbb{BEAST }}\right\}\]

OpenStudy (callisto):

lol I tried to tell you to add ' \ ' as to add a 'space' to separate two words... Better stop here, or we'll bring so much trouble to the asker :|

OpenStudy (anonymous):

LOLOLOL i think we're already in trouble :P